# Reducing base drag on road vehicles using pulsed jets optimized by hybrid genetic algorithms

### v1

@@ 3 Machine learning control
Building on the facility and diagnostics detailed in Section 2, we formulate drag reduction as an open-loop, model-free optimization task over the space of actuation policies [33, 34]. Given the noisy, nonconvex, and potentially discontinuous nature of the objective landscape, we employ a genetically inspired hybrid genetic algorithm that combines global exploration with local refinement, thereby enhancing convergence while exploiting local minima [35]. This section introduces the problem formulation, details the scaling used to construct the cost function, and outlines the algorithmic framework and training process.
@@ 3.1 Formulation of the optimization problem
We introduce a phase shift only between the lateral actuators since relative phase is ill-defined, and thus not a meaningful control variable, when the signals have different frequencies [36, 37]. Flat-back bluff bodies, however, exhibit a strong lateral vortex-shedding mode with a well-defined dominant frequency [38, 5, 6, 7]. By constraining the lateral actuators to share the same frequency $f_{2}$ and duty cycle $DC_{2}$ while varying only their relative phase $\phi$, we explicitly target this mode, for instance, to implement opposition control.
@@
Steady-jet $1.16$ $0.98$ 11 - - - 11 11 11 -
Best $J$ $0.987$ $0.899$ $0.486$ $0.848$ $0.855$ $0.119$ $0.427$ $0.282$ $0.573$ $24$
Best $J_{a}$ $1.005$ $0.875$ $0.716$ $0.614$ $0.761$ $0.58$ $0.686$ $0.524$ $0.669$ $120$
@@
The admissible parameter space is $\tilde{f}_{i}\in[0.085,\,0.935]$, $DC_{i}\in[0.25,\,0.75]$, $\phi\in[0,\,\pi]$. Each parameter’s range is uniformly discretized into $2^{N_{b}}$ values, where the number of bits, $N_{b}$, is specified for each parameter in Table 1. These boundaries are set to concentrate the search around the expected lateral shedding frequency of flat-back bluff bodies, $\tilde{f}_{\text{shed}}\approx 0.11\;\text{–}\;0.20$ [23, 6].
@@ 3.2 Cost function
The objective function must primarily reflect drag reduction. However, directly minimizing a drag-related quantity can bias the optimizer toward control strategies with excessive mass or momentum injection (i.e., high duty cycles), artificially improving performance metrics without necessarily targeting the underlying flow mechanisms responsible for genuine drag mitigation. To prevent this, the cost function includes an explicit penalty on the injected mass flow rate (as discussed by Castellanos et al. [39] for heat transfer), discouraging unphysical solutions and ensuring that the reported improvements result from effective flow control rather than disproportionately strong actuation. Both objectives are combined into a single, dimensionless cost function to allow consistent comparison across all cases:
@@
where $D$ and $\dot{m}$ are the episode-averaged (whose duration is defined in 0.A) drag and mass-flow rate, $D_{0}$ is the mean drag with no actuation, and $\dot{m}_{\mathrm{SJ}}$ is the mean mass flow when all actuators are continuously on (hereafter, the steady-jet reference).
The dimensionless drag term, $J_{a}=D/D_{0}$, quantifies the effectiveness of each candidate actuation (lower values indicate better performance) and depends on all control parameters. The penalization term, $J_{b}=\dot{m}/\dot{m}_{\mathrm{SJ}}$, represents the injected mass flow relative to the steady-jet reference. Under binary actuation with fixed supply conditions, $J_{b}$ depends solely on the duty cycles, being independent of frequency and phase. With this normalization, the no-actuation baseline yields $J=1$. An in-depth description of the evaluation procedure for each individual is provided in Appendix 0.A.
The weight $\gamma\geq 0$ defines the trade-off between drag reduction and mass-flow penalization. Selecting $\gamma$ is non-trivial, as it determines which region of the Pareto front the single-objective optimizer emphasizes. If $\gamma=0$, the optimizer favors solutions that minimize $J_{a}$ at the expense of excessive mass injection, which is energy-inefficient and not necessarily indicative of genuine drag mitigation. Conversely, excessively large $\gamma$ values drive the search toward negligible mass injection, suppressing possible improvements in $J_{a}$. Our goal is to select $\gamma$ such that $J$ is effectively insensitive to variations in mass injection across admissible actuation policies, thereby focusing the optimization on drag reduction.
Following Castellanos et al. [39], we generated $100$ Latin Hypercube samples (LHS) to initialize the optimizer and approximately space-fill the parameter space. Evaluating these samples indicated that $\gamma=0.182$ successfully removes the bias toward large $J_{b}$ without introducing an artificial preference for small $J_{b}$. All subsequent results use this fixed $\gamma$.
@@ 3.3 Hybrid genetic algorithm optimization
Given the complex optimization landscape, featuring nonlinear couplings and collinearity, possible non-smoothness, and a moderately high dimensionality, we employ a hybrid genetic algorithm. The optimizer, labelled HyGO and introduced by Robledo et al. [35], accelerates convergence by appending a local-refinement step to each generation while retaining the global search capability of standard GAs.
An individual encodes the parameter vector in Equation (2). As stated in Section 3.2, the search is initialized via Latin Hypercube Sampling with $N_{init}=100$ individuals. Each subsequent generation proceeds in two stages. First, an explorative stage (generation 1 uses the LHS set as its exploration pool) applies standard GA operators to produce $N_{explor}$ offspring, which are then evaluated and ranked. Second, an exploitative stage generates $N_{exploit}$ additional candidates by locally refining selected high-ranked individuals using the Downhill Simplex method (Nelder–Mead), chosen for its gradient-free and fast-convergent behavior. The number of generations $N_{g}$, tournament size for exploration, probabilities of genetic operators, and other hyperparameters are reported in Table 1 and were selected based on prior experience and related studies [40, 41, 39]. We employ a decreasing exploration population across generations, which front-loads exploration to sample multiple local minima early, and progressively shifts budget toward local refinement.
@@
Figure 5: Time-averaged ($\overline{p}$) and fluctuating ($p^{\prime}$) base pressure contours for four characteristic cases. From left to right, the columns correspond to: (i) the non-actuated baseline flow; (ii) continuous steady-jet actuation; (iii) the optimal individual achieving the minimum total cost, $\min(J)$; and (iv) the individual achieving the maximum drag reduction, $\min(J_{a})$. Each image limits represent entire base surface.
@@ 4.1 Optimization results
A total of $688$ individuals were evaluated over $10$ generations. Thanks to the uncertainty-screening protocol described in the Appendix 0.A, no outliers were retained in the analysis. The hybrid optimizer exhibited fast convergence: it explored the parameter space broadly in the early generations and then converged rapidly in all parameters except $\phi$. Convergence is further supported by the plateau of the best objective in the final three generations (Figure 3(a)), where no further improvement is observed.
@@
A clear optimum region emerges for the three actuation frequencies: $\tilde{f}\approx 0.85$ for the top and side slots ($\tilde{f}_{1},\tilde{f}_{2}$) and $\tilde{f}\approx 0.12$ for the bottom slot, the latter lying close to the natural shedding frequency of the wake. Some scatter remains in $\tilde{f}_{1}$, with a secondary local minimum appearing in later generations near $\tilde{f}_{1}\approx 0.22$. Furthermore, the optimizer tends toward intermediate blowing levels for the top and bottom diffusers while minimizing the side blowing. Since $DC_{2}$ regulates two slots (whereas $DC_{1}$ and $DC_{3}$ each regulate one), the objective $J$ may implicitly favor reducing $DC_{2}$ to limit total mass flow, leaving drag reduction to the other actuators. This remains a hypothesis; targeted sensitivity tests would be required to confirm it.
@@
Interestingly, $J_{a}$ attains low values from the beginning, so HyGO prioritizes maintaining $J_{a}$ at a “good-enough” level while reducing the momentum input (thrust), thereby targeting genuine drag reduction mechanisms, an evolution not immediately apparent in Figure 3(b). The Pareto analysis in Figure 4(b) corroborates this behavior: the best early individuals lie near the minimum of $J_{a}$, and subsequent generations progress downward along the right end of the front, achieving lower $J_{b}$ at nearly constant $J_{a}$. Consistently, Figure 4(a) indicates that the chosen trade-off parameter $\gamma$ was effective: the overall optimum sits at the elbow of the front, delivering substantial drag reduction at an intermediate blowing level. A similar trend has been reported for the same van model when the actuation law is obtained by DRL. Amico et al. [29] trained agents either to maximize drag reduction or to maximize drag reduction under an energy penalty; in the former case, the solution used stronger bottom injection and produced the largest wake compression, whereas in the energy-penalized case, the actuation levels were lower and the wake remained closer to the baseline topology. The herein discussed optimization by HyGO reproduces this behavior (while obtaining better drag reduction, 10% vs 5% under actuation penalization) without changing the cost definition between experiments: early individuals sit near the “maximum-drag-reduction” part of the front (low $J_{a}$, high $J_{b}$), and subsequent generations move towards the energy-efficient branch while keeping $J_{a}$ nearly constant.
@@
Table 2 summarizes the metrics and settings for the overall best individual (minimum $J$), the reference cases (no actuation and steady-jet), and the individual that maximizes drag reduction (minimum $J_{a}$). The steady-jet reference is clearly inefficient: the improvement is limited (about $2\%$), while the net cost remains unfavorable once the mass-injection penalty is included. By contrast, the overall optimum achieves $J<1$, i.e., a net benefit even after penalizing mass injection. The configuration that maximizes drag reduction reaches $\approx 12.5\%$ reduction but does so at very high mass injection, resulting in $J>1$ and therefore worse than no actuation in cost–benefit terms. The global optimum strikes a better balance, delivering $\approx 10\%$ drag reduction while using roughly half of the available mass-flow budget, primarily by acting through the top and bottom slots.
For the optimal control, the bottom slot operates at a low frequency ($\tilde{f}_{3}\approx 0.12$), close to the natural vortex shedding mode ($\tilde{f}_{\mathrm{sh}}\approx 0.17$ in our configuration). We hypothesize that this actuation targets the wake’s global instability, directly manipulating the formation and advection of the large-scale von Kármán vortices [42]. Simultaneously, the top slot provides high-frequency forcing ($\tilde{f}_{1}/\tilde{f}_{3}\approx 7.08$), which may be modulating the shear-layer instabilities separating from the body’s surface. By introducing small-scale vorticity into the shear layers, this high-frequency actuation could disrupt the initial roll-up process that feeds the larger downstream wake structures. While this interpretation is compelling, its validation would require further time-resolved analysis and a consideration of three-dimensional effects that could introduce broadband spectral content.
@@ 4.2 Base pressure analysis
For a more detailed analysis of the flow dynamics, time-resolved pressure data were acquired for two-minute intervals at a sampling frequency of $f_{s}=800$ Hz. The resulting mean ($\overline{\Delta P}=\overline{P-P_{\infty}}$) and fluctuating ($\Delta P^{\prime}=\overline{|\Delta P-\overline{\Delta P}|}$) differential pressure maps for these key scenarios are presented in Figure 5. In the non-actuated case, the mean pressure map shows a pronounced low-pressure region near the lower part of the base. This feature is associated with the dominant shedding mechanism, which produces a large separated region, and the overall pressure distribution is consistent with previous studies [27]. In contrast, all actuation strategies induce significant pressure recovery in this lower area, albeit at the expense of increased pressure losses near the upper edge. The optimal (minimum $J$) and maximum drag reduction (minimum $J_{a}$) cases yield similar mean pressure profiles, characterized by a more extensive area of pressure recovery than that achieved with steady blowing, though with a lower peak pressure, supporting the hypothesis of the actuation targeting the lower separation region.
@@
An analysis of the pressure fluctuations reveals that all control strategies increase the unsteadiness relative to the baseline case. Steady blowing, in particular, generates a large region of high-amplitude fluctuations in the central part of the base, a phenomenon likely attributable to the impingement and shedding of large-scale von Kármán-like vortices. Notably, the pulsed actuation cases produce even higher fluctuation levels, with an asymmetry biased towards the right side of the base. This spatial bias is hypothesized to be a direct consequence of the imposed phase shift between the lateral actuators, wherein the left-side jet systematically lags the right. Furthermore, the greater mass flow rate associated with the maximum drag reduction case (minimum $J_{a}$) correlates with a further increase in fluctuation intensity, including a distinct peak in $\Delta P^{\prime}$ near the bottom of the measurement domain. In general, this spatial organization of the pressure fields is also consistent with the wake topologies reported by Amico et al. [29] for their DRL-controlled van. In their maximum-drag-reduction cases, the streamwise bubble is shortened, the wake becomes more symmetric, and a substantial pressure recovery is measured over the central and lower areas of the base. In our study, the optimal-$J$ and optimal-$J_{a}$ individuals generate the same signature (pressure gain low, mild losses high) but with an actuation found under an explicit mass-flow penalty. This supports our interpretation that HyGO is driving the flow toward the same physically efficient wake state identified by closed-loop DRL, only through an open-loop, model-free search.
Figure 6: Identification of distinct control regimes via clustering. The left image displays the GMM clustering of all evaluated control cases in the reduced-order space ($\alpha_{1},\alpha_{2}$). The six identified clusters are numbered and colored according to their mean cost $J$, from best-performing (Cluster 1) to worst-performing. The right panels include the weighted-average base pressure maps corresponding to each cluster. The top row displays the mean pressure ($\overline{\Delta P}$) and the bottom row shows the pressure fluctuations ($\Delta P^{\prime}$). The maps are ordered by cluster number from left to right, revealing the characteristic flow topology associated with each performance level.
@@
Synchronous pressure and force measurements were acquired for all actuation cases (as detailed in Appendix 0.A). These measurements yielded mean $\overline{\Delta P}$ and fluctuating $\Delta P^{\prime}$ pressure maps for each individual, averaged over the valid repetitions. To facilitate interpretation, the concatenated pressure data were processed using Multi-Dimensional Scaling (MDS) [43, 44], a dimensionality reduction technique. MDS projects the high-dimensional data, 62 dimensions arising from 31 two-component pressure measurements, into a low-dimensional embedding. A correlation analysis between the control parameters and the embedding dimensions reveals that significant information is contained only within the first two dimensions, $\alpha_{1}$ and $\alpha_{2}$, allowing for a visual representation of the explored solution space during the optimization process.
A soft clustering analysis was performed to partition the data into distinct operational regimes within the low-dimensional embedding represented by $\alpha_{1}$ and $\alpha_{2}$. The method utilizes a Gaussian Mixture Model (GMM) [45], a probabilistic model that assigns each data point a posterior probability of belonging to one of several clusters. These probabilities are subsequently localized by retaining assignments only to a point’s most likely cluster and its immediate neighbors, as determined by a k-nearest neighbors (k-NN) search. The resulting six clusters, illustrated in Figure 6, were ordered according to their mean cost function value, $\overline{J}^{k}$, computed as the weighted average of the cost, $J_{i}$, for each individual $i$:
$\overline{J}^{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{ind}}p^{k}_{i}\cdot J_{i},\quad\text{where}\quad N_{k}=\sum_{i=1}^{N_{ind}}p^{k}_{i}.$ (6)
Here, $p^{k}_{i}$ is the localized probability of individual $i$ belonging to cluster $k$, and $N_{ind}$ is the total number of individuals.
@@
Visualizing the principal directions, $\alpha_{1}$ and $\alpha_{2}$, reveals significant trends in the optimization process (Figure 6). The high-performing individuals (i.e., those with lower cost function values) collapse in a specific region of the low-dimensional embedding, providing a clear visual demonstration of the optimization’s convergence. It is particularly noteworthy that this embedding was derived from the pressure measurements, not the actuation parameters themselves. This indicates a smooth mapping between the control inputs and the resulting pressure field, a feature effectively exploited by the algorithm to enhance convergence.
Following the cluster classification, and weighting the pressure maps following the same procedure as in Equation (6), the corresponding weighted-averaged mean and fluctuating pressure maps for each cluster are depicted in Figure 6. The results reveal a clear trend: higher-performing clusters are characterized by a significant pressure increase on the lower half of the base, which is associated with a beneficial modification of the large-scale vortex dynamics, along with minimized pressure fluctuations. Conversely, lower-performing clusters display a pressure peak positioned higher on the base and are accompanied by large pressure oscillations near the lower edge of the measurement domain.

### v2

@@ 3 Machine learning control
Building on the facility and diagnostics detailed in Section 2, we formulate drag reduction as an open-loop, model-free optimization task over the space of actuation policies [BruntonNoack2015review, duriez2017book]. Given the noisy, nonconvex, and potentially discontinuous nature of the objective landscape, we employ a genetically inspired hybrid genetic algorithm that combines global exploration with local refinement, thereby enhancing convergence while exploiting local minima [Robledo2025Hygo]. This section introduces the problem formulation, details the scaling used to construct the cost function, and outlines the algorithmic framework and training process.
@@ 3.1 Formulation of the optimization problem
We introduce a phase shift only between the lateral actuators since relative phase is ill-defined, and thus not a meaningful control variable, when the signals have different frequencies [Wu2018jet, minelli2020lgac]. Flat-back bluff bodies, however, exhibit a strong lateral vortex-shedding mode with a well-defined dominant frequency [evrard2016fluidSymmetryModes, dalla2019simulationsBimodalFlatBackBluff, fan2020experimentalBiStableFlatBackAhmed, khan2024equilibriumFluxesFlatBackBluffBody]. In particular, our model ($H/W\approx 1.17$) is expected to present both lateral and vertical shedding instabilities [Grandemange2013BistabilityBackAR] (but dominated by the vertical one). By constraining the lateral actuators to share the same frequency $f_{2}$ and duty cycle $DC_{2}$ while varying their relative phase $\phi$ we specifically target the lateral, symmetric instability mode, which has been shown [KHAN2022ControlBistable] to contribute significantly to drag when present, and whose partial attenuation, rather than complete suppression, was demonstrated to yield effective drag reduction.
It is also noted that, in the present parametrization, the duty cycle $DC_{2}$ simultaneously controls two actuators, whereas $DC_{1}$ and $DC_{3}$ each affect only a single slot. As a result, variations in $DC_{2}$ have a larger impact on the total injected mass flow and on the global cost function, which implies a higher optimization sensitivity to this parameter and may inherently bias the search toward reducing $DC_{2}$ relative to the other duty cycles.
@@
Steady-jet $1.193$ $1.011$ 11 - - - 11 11 11 -
Best $J$ 11 $0.912$ $0.486$ $0.848$ $0.855$ $0.119$ $0.427$ $0.282$ $0.573$ $24$
Best $J_{a}$ $1.02$ $0.89$ $0.716$ $0.614$ $0.761$ $0.58$ $0.686$ $0.524$ $0.669$ $120$
@@
The admissible parameter space is $\tilde{f}_{i}\in[0.085,\,0.935]$, $DC_{i}\in[0.25,\,0.75]$, $\phi\in[0,\,\pi]$. Each parameter’s range is uniformly discretized into $2^{N_{b}}$ values, where the number of bits, $N_{b}$, is specified for each parameter in Table 1. These boundaries are set to concentrate the search around the expected lateral shedding frequency of flat-back bluff bodies, $\tilde{f}_{\text{shed}}\approx 0.11\;\text{–}\;0.20$ [li2017GP_FlatBackAhmed, fan2020experimentalBiStableFlatBackAhmed].
@@ 3.2 Cost function
The objective function must primarily reflect drag reduction. However, directly minimizing a drag-related quantity can bias the optimizer toward control strategies with excessive mass or momentum injection (i.e., high duty cycles), artificially improving performance metrics without necessarily targeting the underlying flow mechanisms responsible for genuine drag mitigation. To prevent this, the cost function includes an explicit penalty on the injected mass flow rate (as discussed by castellanos2023genetically for heat transfer), discouraging unphysical solutions and ensuring that the reported improvements result from effective flow control rather than disproportionately strong actuation. Both objectives are combined into a single, dimensionless cost function to allow consistent comparison across all cases:
@@
where $D$ and $\dot{m}$ are the episode-averaged (whose duration is defined in A) drag and mass-flow rate, $D_{0}$ is the mean drag with no actuation, and $\dot{m}_{\mathrm{SJ}}$ is the mean mass flow when all actuators are continuously on (hereafter, the steady-jet reference).
The dimensionless drag term, $J_{a}=D/D_{0}$, quantifies the effectiveness of each candidate actuation (lower values indicate better performance) and depends on all control parameters. The penalization term, $J_{b}=\dot{m}/\dot{m}_{\mathrm{SJ}}$, represents the injected mass flow relative to the steady-jet reference. Under binary actuation with fixed supply conditions, $J_{b}$ depends solely on the duty cycles, being independent of frequency and phase. With this normalization, the no-actuation baseline yields $J=1$. An in-depth description of the evaluation procedure for each individual is provided in Appendix A. The weight $\gamma\geq 0$ defines the relative contribution of the mass-injection term $J_{b}$ to the global cost function $J$. In the present formulation, $\gamma$ is not only introduced as a subjective trade-off parameter between aerodynamic performance and actuation cost, but also to decouple genuine wake-induced drag reduction from artificial improvements derived from excessive mass injection. When $\gamma=0$, the optimizer systematically favors solutions with large $J_{b}$, leading to a reduction of $J_{a}$ that is not necessarily associated with beneficial wake modification. Conversely, excessively large values of $\gamma$ drive the optimization toward vanishing actuation, suppressing physically meaningful control strategies. The role of $\gamma$ is therefore set to compensate for this mass-injection effect and to ensure that the ranking of individuals is primarily governed by their ability to modify the wake and recover base pressure. Following castellanos2023genetically, we initialized the optimization using 100 Latin Hypercube samples (that are later taken as the first generation for the optimization) to probe the parameter space and quantify the correlation between $J_{a}$ and $J_{b}$. For $\gamma=0$, these samples exhibit a strong linear dependence between mass injection and apparent drag reduction, indicating that higher blowing levels trivially improve $J_{a}$ through mass injection. The value $\gamma=0.182$ was obtained by linear regression so as to flatten this trend, effectively neutralizing the direct contribution of mass injection to the cost function. Although the initial LHS sampling is necessarily sparse in a seven-dimensional experimental space, the robustness of this choice is validated by the convergence of the optimization toward non-trivial solutions. All subsequent results, therefore, employ this fixed value of $\gamma$.
@@ 3.3 Hybrid genetic algorithm optimization
Given the complex optimization landscape, featuring nonlinear couplings and collinearity, possible non-smoothness, and a moderately high dimensionality, we employ a hybrid genetic algorithm. The optimizer, labelled HyGO and introduced by Robledo2025Hygo includes both GA and GP functionalities but, due to the open-loop and parametric nature of the target problem, the hybridized GA was employed. Furthermore, HyGO accelerates convergence by appending a local-refinement step to each generation while retaining the global search capability of standard GAs.
An individual encodes the parameter vector in Equation (2). As stated in Section 3.2, the search is initialized via Latin Hypercube Sampling with $N_{init}=100$ individuals. Each subsequent generation proceeds in two stages. First, an explorative stage (generation 1 uses the LHS set as its exploration pool) applies standard GA operators to produce $N_{explor}$ offspring, which are then evaluated and ranked. Second, an exploitative stage generates $N_{exploit}$ additional candidates by locally refining selected high-ranked individuals using the Downhill Simplex method (Nelder–Mead), chosen for its gradient-free and fast-convergent behavior. The number of generations $N_{g}$, tournament size for exploration, probabilities of genetic operators, and other hyperparameters are reported in Table 1 and were selected based on prior experience and related studies [Yu2021GA_drag_slot_jets, cornejo2021gMLC, castellanos2023genetically]. We employ a decreasing exploration population across generations, which front-loads exploration to sample multiple local minima early, and progressively shifts budget toward local refinement.
@@
Figure 5: Time-averaged ($\overline{p}$) and fluctuating ($p^{\prime}$) base pressure contours for four characteristic cases. From left to right, the columns correspond to: (i) the non-actuated baseline flow; (ii) continuous steady-jet actuation; (iii) the best-performing individual achieving the minimum total cost, $\min(J)$; and (iv) the individual achieving the maximum drag reduction, $\min(J_{a})$. Each image limits represent entire base surface.
@@ 4.1 Optimization results
A total of $688$ individuals were evaluated over $10$ generations. Thanks to the uncertainty-screening protocol described in the Appendix A, no outliers were retained in the analysis. The hybrid optimizer exhibited fast convergence: it explored the parameter space broadly in the early generations and then converged rapidly in all parameters except $\phi$. Convergence is further supported by the plateau of the best objective in the final three generations (Figure 3(a)), where no further improvement is observed.
@@
A clear optimum region emerges for the three actuation frequencies: $\tilde{f}\approx 0.85$ for the top and side slots ($\tilde{f}_{1},\tilde{f}_{2}$) and $\tilde{f}_{3}\approx 0.12$ for the bottom slot, the latter lying close to the natural shedding frequency of the wake. Some scatter remains in $\tilde{f}_{1}$, with high-performing individuals sparsely distributed across its domain. Furthermore, the optimizer tends toward intermediate blowing levels for the top and bottom diffusers while minimizing the side blowing. Since $DC_{2}$ regulates two slots (whereas $DC_{1}$ and $DC_{3}$ each regulate one), the objective $J$ may implicitly favor reducing $DC_{2}$ to limit total mass flow, leaving drag reduction to the other actuators. This remains a hypothesis; targeted sensitivity tests would be required to confirm it. Consistently with this reduced lateral authority, the best-performing solutions exhibit a large scatter in the lateral phase $\phi$, with both $\phi\approx 24^{\circ}$ and $\phi\approx 120^{\circ}$ yielding comparable performance in the best $J$ and $J_{a}$ solutions, indicating that the objective function is only weakly sensitive to the side-slot phase in the vicinity of the found minimum.
@@
Interestingly, $J_{a}$ attains low values from the beginning, so HyGO prioritizes maintaining $J_{a}$ at a “good-enough” level while reducing the momentum input, thereby targeting genuine drag reduction mechanisms, an evolution not immediately apparent in Figure 3(b). The Pareto analysis in Figure 4(b) corroborates this behavior: the best early individuals lie near the minimum of $J_{a}$, and subsequent generations progress downward along the right end of the front, achieving lower $J_{b}$ at nearly constant $J_{a}$. Consistently, Figure 4(a) indicates that the chosen trade-off parameter $\gamma$ was effective: the overall optimum sits at the elbow of the front, delivering substantial drag reduction at an intermediate blowing level. A similar trend has been reported for the same van model when the actuation law is obtained by DRL. amico2024flow trained agents either to maximize drag reduction or to maximize drag reduction under an energy penalty; in the former case, the solution used stronger bottom injection and produced the largest wake compression, whereas in the energy-penalized case, the actuation levels were lower and the wake remained closer to the baseline topology. The herein discussed optimization by HyGO reproduces this behavior (while obtaining better drag reduction, 8.8% vs 5% under actuation penalization) without changing the cost definition between experiments: early individuals sit near the “maximum-drag-reduction” part of the front (low $J_{a}$, high $J_{b}$), and subsequent generations move towards the energy-efficient branch while keeping $J_{a}$ nearly constant.
@@
Table 2 summarizes the metrics and settings for the overall best individual (minimum $J$), the reference cases (no actuation and steady-jet), and the individual that maximizes drag reduction (minimum $J_{a}$). The steady-jet reference is clearly inefficient with a cost value much higher than unity and a drag increase $J_{a}>1$. By contrast, the overall optimum achieves $J\approx 1$, i.e., a net benefit even after penalizing mass injection. The configuration that maximizes drag reduction reaches $\approx 11\%$ reduction but does so at very high mass injection, resulting in $J>1$ and therefore worse than no actuation in terms of the mass injection–drag reduction balance. The global optimum strikes a better balance, delivering $\approx 8.8\%$ drag reduction while using roughly half of the available mass-flow budget, primarily by acting through the top and bottom slots.
@@ 4.2 Base pressure analysis
For a more detailed analysis of the flow dynamics, time-resolved pressure data were acquired for two-minute intervals at a sampling frequency of $f_{s}=800$ Hz. The resulting mean ($\overline{\Delta P}=\overline{P-P_{\infty}}$) and fluctuating ($\Delta P^{\prime}=\overline{|\Delta P-\overline{\Delta P}|}$) differential pressure maps for these key scenarios are presented in Figure 5. In the non-actuated case, the mean pressure map shows a pronounced low-pressure region near the lower part of the base. This feature is associated with the dominant shedding mechanism, which produces a large separated region, and the overall pressure distribution is consistent with previous studies [amico2022geneticFurgo]. In contrast, all actuation strategies induce significant pressure recovery in this lower area, albeit at the expense of increased pressure losses near the upper edge. The best-performing (minimum $J$) and maximum drag reduction (minimum $J_{a}$) cases yield similar mean pressure profiles, characterized by a more extensive area of pressure recovery than that achieved with steady blowing, though with a lower peak pressure, supporting the hypothesis of the actuation targeting the lower separation region.
@@
An analysis of the pressure fluctuations reveals that all control strategies increase the unsteadiness relative to the baseline case. Steady blowing, in particular, generates a large region of high-amplitude fluctuations in the central part of the base, a phenomenon likely attributable to the impingement and shedding of large-scale von Kármán-like vortices. Notably, the pulsed actuation cases produce even higher fluctuation levels, with an asymmetry biased towards the right side of the base. This spatial bias is hypothesized to be a direct consequence of the imposed phase shift between the lateral actuators, wherein the left-side jet systematically lags the right. Furthermore, the greater mass flow rate associated with the maximum drag reduction case (minimum $J_{a}$) correlates with a further increase in fluctuation intensity, including a distinct peak in $\Delta P^{\prime}$ near the bottom of the measurement domain. In general, this spatial organization of the pressure fields is also consistent with the wake topologies reported by amico2024flow for their DRL-controlled van. In their maximum-drag-reduction cases, the streamwise bubble is shortened, the wake becomes more symmetric, and a substantial pressure recovery is measured over the central and lower areas of the base. In our study, the minimal-$J$ and minimal-$J_{a}$ individuals generate the same signature (pressure gain low, mild losses high) but with an actuation found under an explicit mass-flow penalty. This supports our interpretation that HyGO is driving the flow toward the same physically efficient wake state identified by closed-loop DRL, only through an open-loop, model-free search.
@@
Synchronous pressure and force measurements were acquired for all actuation cases (as detailed in Appendix A). These measurements yielded mean $\overline{\Delta P}$ and fluctuating $\Delta P^{\prime}$ pressure maps for each individual, averaged over the valid repetitions. To facilitate interpretation, the concatenated pressure data were processed using Multi-Dimensional Scaling (MDS) [cox2000mds, kaiser2017MDS], a dimensionality reduction technique. MDS projects the high-dimensional data (i.e., 62 dimensions arising from 31 two-component pressure measurements) into a low-dimensional embedding. A correlation analysis between the control parameters and the embedding dimensions reveals that significant information is contained only within the first two dimensions, $\alpha_{1}$ and $\alpha_{2}$, allowing for a visual representation of the explored solution space during the optimization process. In the following, we stress that the embedding is constructed from pressure measurements, not from the actuation parameters, and therefore provides a flow-response-based representation of the optimization landscape.
To interpret the low-dimensional embedding represented by $\alpha_{1}$ and $\alpha_{2}$, a clustering analysis is performed to distinguish different optimization regions. To ensure robustness, the regions are obtained by performing a K-neighbors clustering on a homogeneously downsampled embedding, which is then mapped to the full-dimensional data. The resulting six groups, illustrated in Figure 6, were ordered according to their mean cost function value, $\overline{J}^{k}$, computed as the average of the cost, $J_{i}$, for each individual $i$ in group $k$:
$\overline{J}^{k}=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}J_{i}$ (6)
where $N_{k}$ is the number of individuals in group $k$. We emphasize that the clustering is introduced primarily for visualization and interpretation: this low-order embedding represents a continuous family of pressure distributions, and the clusters (labelled groups to emphasize the method’s target) serve as a compact discretization into representative regions rather than as strictly separated physical modes.
Figure 6: Identification of distinct control regimes via clustering. The left image displays the clustering of all evaluated control cases in the reduced-order space ($\alpha_{1},\alpha_{2}$). The six identified groups are numbered and colored according to their mean cost $J$, from best-performing (Group 1) to worst-performing. The right panels include the weighted-average base pressure maps corresponding to each group. The top row displays the mean pressure ($\overline{\Delta P}$) and the bottom row shows the pressure fluctuations ($\Delta P^{\prime}$). The maps are ordered by group number from left to right, revealing the characteristic flow topology associated with each performance level.
@@
Visualizing the principal directions, $\alpha_{1}$ and $\alpha_{2}$, reveals significant trends in the optimization process (Figure 6). The high-performing individuals (i.e., those with lower cost function values) concentrate within a compact region of the low-dimensional embedding, providing a clear visualization of the optimization’s convergence. As previously remarked, it is particularly noteworthy that this embedding was derived from the pressure measurements, not the actuation parameters themselves. This indicates a smooth relationship within the explored domain between the control inputs and the resulting pressure field, a feature effectively exploited by the algorithm to enhance convergence. Furthermore, the relative locations of the best $J$ and best $J_{a}$ solutions within the embedding indicate that the pressure manifold separates flow regimes not only according to aerodynamic performance, but also according to actuation-cost efficiency. The trade-off between $J_{a}$ and $J_{b}$ is also evident from the group-wise statistics: the fourth-best group according to $J$ is ranked third when evaluated in terms of $J_{a}$, indicating that this group achieves strong drag reduction at the expense of significantly higher mass injection. This embedding therefore reveals a preferential organization of solutions, with a predominantly vertical stratification associated with $J$, where lower regions correspond to actuation–cost performance, and a left–right organization correlated with $J_{a}$, reflecting better aerodynamic performance.
Following this classification, and averaging the pressure maps following the same procedure as in Equation (6), the corresponding average mean and fluctuating pressure maps for each group are depicted in Figure 6. The results reveal a clear trend: higher drag-reduction groups are characterized by stronger mean pressure recovery concentrated over the lower portion of the base, together with comparatively reduced pressure fluctuations. Progressing from group 1 to group 6, the fluctuation levels $\Delta P^{\prime}$ increase markedly, with the most intense fluctuations localized near the lower base region in the low-performance regimes. In parallel, the mean-pressure footprint becomes less favorable on the lower base and appears increasingly localized and/or shifted upward in the worst-performing groups. An exception to this trend is observed in the fourth group, which exhibits a $\overline{\Delta P}$ distribution similar to that of the best-performing group due to its large drag reduction (low $J_{a}$); however, the higher mass injection leads to increased pressure fluctuations (as evidenced in the $\Delta P^{\prime}$ map), placing this group in the fourth position in terms of $J$. Overall, these group-wise pressure signatures provide a compact overview of the wake regimes explored by the optimizer and their association with aerodynamic performance.